OpenStudy (anonymous):

The graph of a parabola is represented by the equation y=ax^2 where a is a positive integer. If a is multiplied by 2, the new parabola would become? 1. Narrower and open downward 2. Narrower and open upward 3.wider and open downward 4. Open and open upward

OpenStudy (austinl):

It become narrower, and opens upward.

OpenStudy (anonymous):

I think it might be answer #2. But I dont know how to explain/show my work for it.

OpenStudy (anonymous):

How though? That's what i'm mainly stuck on.

OpenStudy (austinl):

Let's work it out! x y=x^2 y = 2x^2 -3 9 18 -2 4 8 -1 1 2 0 0 0 1 1 2 2 4 8 3 9 18 If you graph those points, do you agree with your answer?

OpenStudy (anonymous):

How would you graph those points? ex:(-3,9) like that?

OpenStudy (austinl):

Correct.

OpenStudy (anonymous):

And yeah, i saw that same answer on a previous question exactly like mine, thats why i re asked the question.

OpenStudy (anonymous):

I thought he would have replied... (surithayer)

OpenStudy (austinl):

idk... well thanks for the medal and keep up the good work!

OpenStudy (anonymous):

\[ax ^{2}=y,x ^{2}=\frac{ 1 }{ a }y\] it is an upward parabola ,because a is positive ,so1/a is also positive when a is multiplied by 2,then 2a>a and1/2a <1/a so length of latus rectum decreases hence it becomes narrower.

OpenStudy (anonymous):

Aww wow, thanks buddy!

OpenStudy (anonymous):

yw